[Hallc_running] [New Logentry] RC Daily Update

dhamilto at jlab.org dhamilto at jlab.org
Sat Dec 1 15:45:01 EST 2018

Logentry Text:

- A very settled period of data-taking since yesterday.
- In order to understand various rate-dependent effects, we are taking two runs at 16uA, then two at 24 uA followed by two at 32uA. A decision will be made soon to choose the optimal current.
- The owl crew were able to see a lambda peak in missing mass data after a process of cuts optimisation, and yields are in line with what was expected.
- Mark looked at the issue of electronic dead time reporting and noticed a factor of 100 missing for the OG 6 GeV reports. The corrected numbers are in line with what is expected (50e-9 [s] * SHMS 3/4 rate [/s]). The electronic deadtime at 16 uA was around 1.1%, while at 35 uA (yesterday) it was 2.7% (see https://logbooks.jlab.org/entry/3633741).
- BCM4C is missing in the SHMS scalers, but is present in the HMS ones. On a related note, Dave G is concerned that the Unser is not reading out in either set of scalers (see https://logbooks.jlab.org/entry/3633696).

RC Meeting:
- There will be no RC meeting today, although I will be around the counting house for the next few hours if anyone wants to discuss anything.

Notices in effect for shift crews:
- Shift leaders, please fill out the check list once per shift.
- Shift leaders, please do a quick cross-check of the BTA numbers against your shift summary.
- Check that the pre-scaler gui is active and hasn't frozen / doesn't need re-started.
- Be on the lookout for any errors when running the replay, particularly errors concerning missing reference times.
- Note the new beam positions for this energy (on the whiteboard / note on wall)
- NB If the HV trips for SHMS S1Y5B contact Simona M. '''DO NOT change the HV from the nominal value!'''

Planned accelerator configuration changes:
- Beam studies / system recovery planned for Monday, Day shift.

Planned accesses to the hall on Monday (during beam studies): 
- Mark and Ethan plan to reboot the HMS collimator motion controller.
- Joe will look at possibly changing the A/C filter on the HMS shield hut..


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3633793
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181201/672d556f/attachment.html>

More information about the Hallc_running mailing list